OpenStudy (anonymous):

Write the complex number in rectangular form: 4(cos(11π/6)+isin(11π/6))

OpenStudy (loser66):

\[cos \frac {11\pi}{6}= \frac{\sqrt{3}}{2} \\sin\frac{11\pi}{6}=\frac{1}{2}\] therefore, yours is \[4(\frac{\sqrt{3}}{2}, \frac{1}{2})=(2\sqrt{3}, 2)\] Is it clear?

OpenStudy (loser66):

hehe, forgot i at sin

OpenStudy (loser66):

it should be (2 sqrt(3), 2i)

OpenStudy (anonymous):

yes! very clear! thank you so much!

OpenStudy (anonymous):

so it is 2 sqrt(3)+2i right? for rectangular form?

OpenStudy (loser66):

i think so

OpenStudy (anonymous):

thanks!

OpenStudy (loser66):

yw